OpenStudy (anonymous):

What affect does h = 2 have on the parent function y=1/x A Graph moves 2 units to the left B Graph moves 2 units to the right C Graph moves 2 units down D Graph moves 2 units up

OpenStudy (anonymous):

what is \(h\)?

OpenStudy (anonymous):

\[\frac{2}{x}\] \[\frac{1}{x+2}\] \[\frac{1}{x}+2\] ? not clear from the question